# On a Conjecture about Ron Graham’s Sequence

### v1

@@ 1.1 Square-product and corresponding sequences
A square-product sequence or a sequence with the square-product property is an increasing sequence $a_{1}<a_{2}<\dots<a_{t}$ such that $a_{1}a_{2}\cdots a_{t}$ is a square number.
@@ 1.2 An upper bound
Let $a_{1}<a_{2}<\dots<a_{t}$ corresponding sequence to $g(p)$ so that $a_{1}=p$. Since $p\mid a_{1}a_{2}\dots a_{t}$ and the sequence has the square-product property, $p^{2}$ must divide the product. Thus the sequence must contain another term that has $p$ as a factor. Since $2p$ is the smallest integer greater than $p$ that has $p$ as a factor, the corresponding sequence must contain a term that is greater than or equal to $2p$. Thus $a_{t}\geq 2p$, and so $g(p)\geq 2p$.
@@ 1.3 Injectivity of Ron Graham’s Sequence
$n^{\prime}=a^{\prime}_{1}<a^{\prime}_{2}<\dots<a^{\prime}_{t^{\prime}}=g(n^{\prime})$
such that $a_{1}a_{2}\dots a_{t}$ and $a^{\prime}_{1}a^{\prime}_{2}\dots a^{\prime}_{t^{\prime}}$ are both square, and thus $a_{1}a_{2}\dots a_{t}a^{\prime}_{1}a^{\prime}_{2}\dots a^{\prime}_{t^{\prime}}$ is square.
@@ 1.4 Surjectivity of Ron Graham’s Sequence
When $n=0$, $g(0)=0\in(\mathbb{N}_{\geq 0}\setminus\mathbb{P})$. For $n>0$ and a square-product sequence $n=a_{1}<a_{2}<\dots<a_{t}$ we claim that $a_{t}$ cannot be prime. If it were, then $p^{2}\mid a_{1}a_{2}\cdots a_{t-1}p$ and thus $p\mid a_{1}a_{2}\cdots a_{t-1}$. However, this cannot occur since $a_{i}<p$ for all $i<t$. ∎
@@
Now we need to check that $g(\bar{g}(k))$ is not strictly less than $k$. For the sake of contradiction, suppose it were—that is, there exists a square-product sequence $\bar{g}(k)=a^{\prime}_{1}<a^{\prime}_{2}<\dots<a^{\prime}_{t^{\prime}}<k$. Since all of terms are positive, we can apply the same argument from the proof of Lemma 8 and see that removing the terms that appear in both sequences results in a sequence that starts with a number greater than $\bar{g}(k)$, ends with $k$, and whose product is square. However, this contradicts the definition of $\bar{g}(k)$, and so $g(\bar{g}(k))=k$ whenever $k$ is non-prime. ∎
@@ 2.1 An algorithm to compute $g(n)$
The sequence $a_{1}<a_{2}<\dots<a_{n}$ is a square-product sequence if and only if $\vec{v}(a_{1})+\vec{v}(a_{2})+\dots+\vec{v}(a_{n})=\vec{0}$.
@@
[$\ \vec{v}(9)&\vec{v}(10)&\dots&\vec{v}(15)\\ \vline&\vline&&\vline\end{array}\right]\vec{b}=\vec{v}(8)$
@@ 2.2 Counting corresponding sequences
Let $n=mr^{2}$ where $m$ is the squarefree part of $n$. Then $g(n)=m(r+1)^{2}$.
@@ 4.1 Conjectures
Recall that Example 27 gives examples where $T(n)=t$ for all integers $t\leq 14$, besides $t=2$. We have checked by computer that $T(n)\geq 15$ for all $n\leq 10000$.
@@ 4.2 Related questions and generalizations
$n=a_{1}\leq a_{2}\leq\dots\leq a_{t}=k,$
such that $a_{1}a_{2}\dots a_{t}=N^{m}$ for some integer $N$, and no integer appears in the sequence $m$ or more times.
@@
Let $T_{m}(n)$ and $T^{\prime}_{m}(n)$ be the minimum number of terms and the minimum number of distinct terms in the solution sequence of $g_{m}(n)$
@@
Does the set of solution sequences of $A300516(n)$ have any notable properties?
@@ Acknowledgments
We would like to thank Alec Jones and Art Benjamin for their suggestions and feedback. We would also like to extend a special thank you to Elise Lockwood—the first step of the journey only happened because you braved a trip to Oregon State University’s campus in the snow in December 2013.
We acknowledge that this paper would likely not have been written without contributions from the late Reinhard Zumkeller.

### v2

@@ 1.1 Square-product and corresponding sequences
A square-product sequence or a sequence with the square-product property is an increasing sequence $a_{1}<a_{2}<\cdots<a_{t}$ such that $a_{1}a_{2}\cdots a_{t}$ is a square number.
@@ 1.2 An upper bound
Let $a_{1}<a_{2}<\cdots<a_{t}$ be a corresponding sequence to $g(p)$ so that $a_{1}=p$. Since $p\mid a_{1}a_{2}\cdots a_{t}$ and the sequence has the square-product property, $p^{2}$ must divide the product. Thus the sequence must contain another term that has $p$ as a factor. Since $2p$ is the smallest integer greater than $p$ that has $p$ as a factor, the corresponding sequence must contain a term that is greater than or equal to $2p$. Thus $a_{t}\geq 2p$, and so $g(p)\geq 2p$.
@@ 1.3 Injectivity of Ron Graham’s Sequence
$n^{\prime}=a^{\prime}_{1}<a^{\prime}_{2}<\cdots<a^{\prime}_{t^{\prime}}=g(n^{\prime})$
such that $a_{1}a_{2}\cdots a_{t}$ and $a^{\prime}_{1}a^{\prime}_{2}\cdots a^{\prime}_{t^{\prime}}$ are both square, and thus $a_{1}a_{2}\cdots a_{t}a^{\prime}_{1}a^{\prime}_{2}\cdots a^{\prime}_{t^{\prime}}$ is square.
@@ 1.4 Surjectivity of Ron Graham’s Sequence
When $n=0$, $g(0)=0\in(\mathbb{N}_{\geq 0}\setminus\mathbb{P})$. For $n>0$ and a square-product sequence $n=a_{1}<a_{2}<\cdots<a_{t}$ we claim that $a_{t}$ cannot be prime. If it were, then $p^{2}\mid a_{1}a_{2}\cdots a_{t-1}p$ and thus $p\mid a_{1}a_{2}\cdots a_{t-1}$. However, this cannot occur since $a_{i}<p$ for all $i<t$. ∎
@@
Now we need to check that $g(\bar{g}(k))$ is not strictly less than $k$. For the sake of contradiction, suppose it were—that is, there exists a square-product sequence $\bar{g}(k)=a^{\prime}_{1}<a^{\prime}_{2}<\cdots<a^{\prime}_{t^{\prime}}<k$. Since all of terms are positive, we can apply the same argument from the proof of Lemma 8 and see that removing the terms that appear in both sequences results in a sequence that starts with a number greater than $\bar{g}(k)$, ends with $k$, and whose product is square. However, this contradicts the definition of $\bar{g}(k)$, and so $g(\bar{g}(k))=k$ whenever $k$ is non-prime. ∎
@@ 2.1 An algorithm to compute $g(n)$
The sequence $a_{1}<a_{2}<\cdots<a_{n}$ is a square-product sequence if and only if $\vec{v}(a_{1})+\vec{v}(a_{2})+\cdots+\vec{v}(a_{n})=\vec{0}$.
@@
[$\ \vec{v}(9)&\vec{v}(10)&\cdots&\vec{v}(15)\\ \vline&\vline&&\vline\end{array}\right]\vec{b}=\vec{v}(8)$
@@ 2.2 Counting corresponding sequences
Let $n=mr^{2}$ where $m$ is the squarefree part of $n$. Then $f(n)=m(r+1)^{2}$.
@@ 4.1 Conjectures
Recall that Example 27 gives examples where $T(n)=t$ for all integers $t\leq 14$, besides $t=2$. We have checked by computer that $T(n)<15$ for all $n\leq 10000$.
@@ 4.2 Related questions and generalizations
$n=a_{1}\leq a_{2}\leq\cdots\leq a_{t}=k,$
such that $a_{1}a_{2}\cdots a_{t}=N^{m}$ for some integer $N$, and no integer appears in the sequence $m$ or more times.
@@
Let $T_{m}(n)$ and $T^{\prime}_{m}(n)$ be the minimum number of terms and the minimum number of distinct terms over all corresponding sequences for $g_{m}(n)$
@@
Does the set of corresponding sequences for $A300516(n)$ have any notable properties?
@@ Acknowledgments
We would like to thank Alec Jones, Art Benjamin, and Sarosh Adenwalla for their suggestions, feedback, and corrections. We acknowledge that this paper would likely not have been written without contributions from the late Reinhard Zumkeller. We would also like to extend a special thank you to Elise Lockwood—the first step of the journey only happened because you braved a trip to Oregon State University’s campus in the snow in December 2013.